Estimate the value of a loyal customer of a loyal Volvo automobile owner. Assume the contribution margin is 0.31, the purchase price is $70,000, the repurchase frequency is every four years, and the customer defection rate is 22 percent. Do not round intermediate calculations. Round your answer to the nearest dollar.
Answer #1

Answer:


Estimate value of a loyal customer (VLC)

The value of a loyal customer is the loyalty shown by the regular customers towards a company.

VLC = P * RF * CM * BLC

 P is revenue per unit,

 RF is repurchase frequency,

 CM is contribution margin,

BLC is buyer's life cycle

BLC = 1 / defection rate

Given

P = $70000

RF = 1 / 4

CM = 0.31

BLC = 1 / 0.22

VLC = $70000* (1 / 4) * 0.31 * (1 / 0.22) = $24,659.09

Value of a loyal customer of a loyal Volvo automobile owner is $24,659.09


answered by: Subrahmanyam golla